OpenStudy (ksaimouli):

3x^2-2x+7=0, then (x-1/3)^2

geerky42 (geerky42):

For x²-2x+7=0, use quadratic formula. For (x-1/3)², it equals (x-1/3)(x-1/3) then expand it using FOIL or box method.

OpenStudy (ksaimouli):

those r not two seperate questions those r 1

OpenStudy (ksaimouli):

\[3x^2-2x+7=0\] then \[{ (\frac{ x-1 }{ 3 })}^2\]

geerky42 (geerky42):

You mean, find the value of \(\left(\dfrac{x-1}{3}\right)^2\)?

OpenStudy (ksaimouli):

yes

geerky42 (geerky42):

Well, first, find value of x from x²-2x+7=0. Use quadratic formula. Then plug in value of x into \(\left(\dfrac{x-1}{3}\right)^2\).

OpenStudy (ksaimouli):

non calculator i did that it takes time

geerky42 (geerky42):

Well, it does. Be patience!

geerky42 (geerky42):

I'm surprised you know calculus but not this.

OpenStudy (ksaimouli):

hmm xD i did this figuring out any new methods

OpenStudy (ksaimouli):

as simply as possible